OpenStudy (anonymous):

Which of the following expressions is the conjugate of a complex number with 2 as the real part and 3i as the imaginary part? 2 + 3i 2 − 3i 3i + 2 3i − 2

Nnesha (nnesha):

you just need to change the sign of imaginary number

OpenStudy (anonymous):

here check the answer http://www.agor.io/

OpenStudy (cbarredo1):

Don't do it

Nnesha (nnesha):

what?

OpenStudy (cbarredo1):

Don't click the link, he thinks he is funny 7-7

OpenStudy (cbarredo1):

3i+2

OpenStudy (anonymous):

lol lameeeee but seriously i need help XD

OpenStudy (anonymous):

and ive been asking if it's 3i+2 and you are the only one who answered back. so that's the answer??

OpenStudy (anonymous):

it's 2-3i

Nnesha (nnesha):

read my first comment

OpenStudy (anonymous):

David was looking at his limousine bill and saw that the total number of miles he has ridden in the car was tracked: Month Miles 5 580 6 695 7 810 8 925 The function representing David's miles per month is f(m) = 115m + 5. What does the 115 represent?

OpenStudy (anonymous):

what about this?

Nnesha (nnesha):

what about the first one ?

OpenStudy (anonymous):

2-3i?

OpenStudy (anonymous):

@Nnesha @CBARREDO1 @AlexandruGuja

Nnesha (nnesha):

looks right

OpenStudy (anonymous):

what abut the one i posted as well

OpenStudy (cbarredo1):

115 is the rate isn't it and 5 is the number of miles at start?

OpenStudy (anonymous):

its asking what 115 represents

Nnesha (nnesha):

y =mx+b where m is slope which represent = rate of change or how things change

OpenStudy (anonymous):

so wouldn't it be the miles at the start?

Nnesha (nnesha):

y = mx + b where b is y-intercept which represent starting amount

Nnesha (nnesha):

f(m) = 115m + 5 in this equation b is 5

OpenStudy (anonymous):

im so confused tbh

Nnesha (nnesha):

why?

OpenStudy (anonymous):

nvm just skip th one i posted. i have a different on

OpenStudy (anonymous):

What is the simplified form of the quantity 15 x y squared over x squared minus 5x plus 6 all over the quantity 5 x squared y over 2x squared minus 7x plus 3 ?

OpenStudy (anonymous):

I am hilarious..

OpenStudy (anonymous):

no ur not xD

Nnesha (nnesha):

\[\huge\rm \frac{ \frac{ 15xy^2 }{ x^2-5x+6} }{ \frac{ 5x^2y }{ 2x^2-7x +3} }\] like this ?

OpenStudy (anonymous):

@scenemunster You know you love my presence cx

OpenStudy (anonymous):

yup and totally lol

OpenStudy (anonymous):

lol

Nnesha (nnesha):

first of all change division to multiplication multiply the first fraction by reciprocal of the denominator fraction example \[\huge\rm \frac{ \frac{ a }{ b } }{ \frac{ c }{ d } } = \frac{ a }{ b } \times \frac{ d }{ c}\]

OpenStudy (anonymous):

what goes where?

OpenStudy (anonymous):

@Nnesha ?

Nnesha (nnesha):

\[\huge\rm \frac{\color{reD}{ \frac{ 15xy^2 }{ x^2-5x+6}} }{\color{blue}{ \frac{ 5x^2y }{ 2x^2-7x +3}} }\] red=first fraction a/b blue= 2nd fraction c/d

OpenStudy (anonymous):

im so confused